I have tried many times i need serious help!this equation isn't working out right i don't know why, help!

Solve the linear system by using substitution:
2x+2y=6 and
-3x+5y=(-33)

Answer by mananth(16946)

2x+2y=6
/2
x+y=3
x= 3-y
..
.
-3x+5y=(-33)
-3(3-y)+5y=-33
-9+3y+5y=-33
8y=-24
/8
y=-3
Plug y
in x+y=3
we get x=6
(6,-3)